the length of a rectangular floor is 5 feet less than twice its width the area of the floor is 150 square feet what is width of the room?
Answer by checkley77(12844)

L=2W-5
LW=150
(2W+5)W=150
2W^2+5W-150=0
(2W-15)(W+10)=0
2W-15=0
2W=15
W=7.5 ANS. FOR THE WIDTH.
L=2*7.5+5
L=15+5
L=20 ANS. FOR THE LENGTH.
PROOF:
20*7.5=150
150=150